# Plasma acceleration of polarized particle beams

### v1

@@ 6 Summary and future prospects
Over the last few years, the area of laser-plasma based acceleration methods for spin-polarized particle beams has significantly evolved. This is in parts due to the advances in general plasma-based accelerator research as well as the recent studies on the generation of polarized particle sources.
Up to now, most of the proposed mechanisms rely on pre-polarized targets, which come with certain restrictions on the maximum attainable density and size. Only a few results suggest the utilization of in-situ setups as in the case of polarized electron beams [43, 44]. In the regime of QED physics, the methods rely on radiative polarization [20, 23]. For pre-polarized targets, it has become obvious in most of the theoretical studies that the injection phase of the particles into the accelerating structure is the most crucial for the preservation of polarization. The subsequent acceleration phase, where $\gamma\gg 1$ was shown to only marginally contribute to depolarization. The recent proof-of-principle results with 3He give hope that laser-plasma based polarized targets are a feasible approach [32].
@@
If upcoming experiments further confirm the feasibility of laser-plasma based sources of polarized particles, they will be of significant interest to the community of high-energy physics for deep-inelastic scattering experiments or the fusion community in the scope of polarized fusion for improved cross sections.
Z.G. would like to thank Karen Z. Hatsagortsyan, Christoph H. Keitel, Matteo Tamburini, and Xiaofei Shen for fruitful discussions. This work has been funded in parts by the DFG (projects PU 213/9-1 and BU 2227/5-1) and BMBF (project 05P24PF1). The work of M.B. has been carried out in the framework of the JuSPARC (Jülich Short-Pulse Particle and Radiation Center [142]).

### v2

@@ 6 Summary and future prospects
Over the last few years, the area of laser-plasma based acceleration methods for spin-polarized particle beams has significantly evolved. This is in part due to the advances in general plasma-based accelerator research as well as the recent studies into the generation of polarized particle sources.
Up to now, most of the proposed mechanisms rely on pre-polarized targets, which come with certain restrictions on the maximum attainable density and size. Only a few results suggest the utilization of in-situ setups as in the case of polarized electron beams [53, 54]. In the regime of QED physics, the methods rely on radiative polarization [28, 31]. For pre-polarized targets, it has become obvious in most of the theoretical studies that the injection phase of the particles into the accelerating structure is the most crucial for the preservation of polarization. The subsequent acceleration phase, where $\gamma\gg 1$, was shown to only marginally contribute to depolarization. The recent proof-of-principle results with 3He give hope that laser-plasma based polarized targets are a feasible approach [41].
@@
Laser–plasma acceleration naturally produces ultrashort pulses (on the order of femtoseconds or picoseconds) with extremely high peak currents in the kiloampere range. These values exceed conventional polarized sources, such as GaAs photocathodes, by three to four orders of magnitude in peak current. The combination of femtosecond/picosecond duration and kiloampere currents provides new opportunities for nuclear and particle physics.
For instance, polarized deep inelastic scattering (DIS) experiments have historically been limited by the achievable beam current from conventional polarized injectors [188, 189]. A kiloampere-class polarized beam would enable instantaneous luminosities orders of magnitude higher, significantly improving statistical precision in measurements of the nucleon spin structure. In polarized proton–deuteron or proton–nucleus scattering, high-current polarized injectors would allow precise mapping of spin-dependent nuclear forces [6]. The short pulse duration additionally provides time-resolved probes of nuclear excitations, something not possible with CW polarized beams.
Laser-driven polarized beams may also impact fusion research, where spin alignment can enhance nuclear reaction cross sections. In particular, in the context of proton-boron (p-11B) fusion, polarized protons accelerated in a laser-plasma environment could modify the angular distribution of alpha particles. The reaction yield can be enhanced if the target is also polarized. The ultrashort pulse duration ensures synchronization with compression or ignition stages, potentially enabling polarized fast-ignition schemes in laser-driven fusion systems.
Perhaps the most immediate application of laser-driven polarized particles is as injectors into conventional accelerator facilities. Conventional polarized sources, such as strained-GaAs photocathodes, are constrained to average currents in the milliampere regime. In contrast, laser-plasma injectors can deliver kiloampere-scale polarized bunches that can be captured and accelerated further in RF linacs or storage rings. This approach provides access to high-energy, high-brightness polarized beams well beyond the limits of conventional injectors, supporting future polarized electron-ion colliders (EICs) at higher instantaneous luminosity for spin-structure studies.
Z.G. would like to thank Karen Z. Hatsagortsyan, Christoph H. Keitel, Matteo Tamburini, and Xiaofei Shen for fruitful discussions. L. J. acknowledges the support by the National Science Foundation of China (No. 12388102) and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0890000). This work has been funded in parts by the DFG (projects PU 213/9-1 and BU 2227/5-1) and BMBF (project 05P24PF1). The work of M.B. has been carried out in the framework of the JuSPARC (Jülich Short-Pulse Particle and Radiation Center [190]).